Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the company’s products, a football helmet for the North American market, requires a special plastic. During the quarter ending June 30, the company manufactured 3,100 helmets, using 1,705 kilograms of plastic. The plastic cost the company $12,958.

According to the standard cost card, each helmet should require 0.50 kilograms of plastic, at a cost of $8.00 per kilogram.

Required:

1. What is the standard quantity of kilograms of plastic (SQ) that is allowed to make 3,100 helmets?

2. What is the standard materials cost allowed (SQ × SP) to make 3,100 helmets?

3. What is the materials spending variance?

4. What is the materials price variance and the materials quantity variance?
Answer #1

Standard Quantity allowed = 3100*0.50

= 1550

Standard cost allowed = 1550*8

= 12,400

Materials spending variance = 12,400 - 12,958

= 558 U

Materials Price Variance = (SP-AP) *AQ

= (8 - 12958/1705)*1705

= 682 F

Materials Quantity Variance = (SQ-AQ) *SP

= (1550 - 1705)*8

= 1340 U